33 cows can graze a field in 12 days.How many cows will graze the same field in 9 days?
Answer by ptaylor(2198)

Let x=number of cows needed to graze the field in 9 days
If 33 cows can graze a field in 12 days, then one cow can graze the field in 33*12=396 days. This means that one cow grazes at the rate of 1/396 fields per day;33 cows graze at the rate of 33/396=1/12 field per day
x cows graze at the rate of x/396 fields per day and we are told that this equals 1/9 fields per day; soooooo
x/396=1/9
9x=396
x=44 cows----Number of cows needed to graze the field in 9 days.
Another way to look at it:
It takes 33*12=396 cow-days (as opposed to man-hours) to graze the field.
This means that 1 cow can graze the field in 396/1=396 days
44 cows can graze the field in 396/44=9 days
33 cows can graze the field in 396/33=12 days
etc.
Hope this helps---ptaylor